# Toponium: the smallest bound state and simplest hadron in quantum mechanics

### v1

@@ A.3 Experimental simulation
In this study, MC simulations of $e^{+}e^{-}$ collisions have been performed. The physics events involving the physics process with or without $J_{t}$ are generated using $\tt Whizard$ [109], where the effects of ISR and energy spread have been taken into account. Subsequently, the generated parton-level events from $\tt Whizard$ are passed to $\tt Pythia6$ [122] for parton showering, hadronization, and the decay processes of the short-life hadrons, ensuring a realistic simulation of the final state particles. Finally, $\tt Delphes$ [123], configured with the CEPC detector conditions [12], is used to simulate the detector response, allowing for the reconstruction of observable quantities from the theoretical predictions.
A.3.1 Event generation in pseudo-experiments
The best option to identify $J_{t}$ is in the final state with double $b$-jets. However, jets originating from other quark flavors can sometimes be misidentified as $b$-jets [124] with the $b$-tagging algorithms [125], thus the contribution of other final states ($u\bar{u}$,${}d\bar{d}$$s\bar{s}$$c\bar{c}$) should be also generated. Moreover, due to limitations in detector performance, some jets with low energy or those produced at extreme angles relative to the detector’s central axis [12, 14] may go undetected, which results in an underestimation of the jet multiplicity in some events. To address this, several multi-jet backgrounds from $WW$, $ZZ$, and $ZH$ decays to four-quark final states are simulated. In order to enhance the sensitivity for $J_{t}$ signal, the event selection criteria are optimized as discussed below.
A.3.2 Event selection criteria
Jets from $q\bar{q}$ are clustered from all the well reconstructed tracks with an anti-kt algorithm [126]. To select high-quality jets, we require jet transverse momentum ($p_{t}$) greater than 15 GeV. To include the case of gluon final state radiation in $q\bar{q}$ final state, the selected events are supposed to have 2 or 3 jets. As the detector does not cover all directions, the ISR photons and jets close to the beam direction may escape detection. Thus, the effective invariant mass of the final states $\sqrt{s^{\prime}}$ needs to be considered. Consequently, $\sqrt{s^{\prime}}$ in Eq.(16) can not be obtained directly by the energies of jets [44], but can be measured well by using the information of jet angular detection with an accuracy of $0.01\%$ [127]. Based on the conservation of four-momentum of the initial and final states, $\sqrt{s^{\prime}}$ of the three-jet final states can be determined as:
$\sqrt{s^{\prime}}=\begin{cases}\sqrt{s}-\sqrt{s}\frac{\cos\theta_{1}\sin\phi_{23}+\sin\theta_{1}\left(\cot\theta_{2}\sin\phi_{31}-\cot\theta_{3}\sin\phi_{21}\right)}{\left(\cos\theta_{1}+1\right)\sin\phi_{23}+\sin\theta_{1}\left(\cot\frac{\theta_{2}}{2}\sin\phi_{31}-\cot\frac{\theta_{3}}{2}\sin\phi_{21}\right)}&\text{(Forward)}\\ \sqrt{s}-\sqrt{s}\frac{\cos\theta_{1}\sin\phi_{23}+\sin\theta_{1}\left(\cot\theta_{2}\sin\phi_{31}-\cot\theta_{3}\sin\phi_{21}\right)}{\left(\cos\theta_{1}-1\right)\sin\phi_{23}+\sin\theta_{1}\left(\tan\frac{\theta_{3}}{2}\sin\phi_{21}-\tan\frac{\theta_{2}}{2}\sin\phi_{31}\right)}&\text{(Backward)}\end{cases},$ (24)
where $\theta_{i}$ and $\phi_{i}$ are the polar and azimuthal angles of the $i-$th jet for $i=1,~{}2,~{}3$, and $\phi_{ij}=\phi_{i}-\phi_{j}$. Forward (backward) indicates that the momentum direction of the ISR photon or missing jet is the same as (opposite to) the direction of the electron beam. In the two cases of forward and backward directions, $1-{\sqrt{s^{\prime}}}/{\sqrt{s}}$ is positive in one case and negative in the other, where the positive value represents the physical solution. Similarly, in the two-jet final states, we can get:
$s^{\prime}=\frac{\sin{\theta_{1}}+\sin{\theta_{2}}-|\sin{(\theta_{1}+\theta_{2})}|}{\sin{\theta_{1}}+\sin{\theta_{2}}+|\sin{(\theta_{1}+\theta_{2})}|}\times s,$ (25)
which is consistent with the Eq.(1) of Ref. [44]. We require $\sqrt{s^{\prime}}>\sqrt{s}-10~{}{\rm GeV}$ here [128, 44, 63].
@@ A.3.2 Event selection criteria
For the measurement of $R_{b}$, only events with at least two $b$-tagged jets [125] are categorized as $b\bar{b}$ events, while the remaining events are considered as light quark pair processes. For the selected signal events, we require the invariant mass of jet pairs $M_{bb}$ or $M_{j_{1}j_{2}}$ greater than 150 GeV to suppress the backgrounds from W/Z hadronic decays, where ${j_{1}j_{2}}$ denote the two jets with the highest and the second-highest energies. Events passing all of the criteria are recognized as signal events, and the detection efficiencies of different processes in the three energy points are shown in Table 3.
@@
Table 3: The experimental cross sections $\sigma(\sqrt{s})$ in fb, double $b$ tagged efficiencies $\varepsilon_{\text{tag}}(\%)$, and detection efficiencies $\varepsilon_{\text{hadron}}(\%)$ for different processes in $R_{b}$ measurements. Here $jj$ is $u\bar{u}$,${}d\bar{d}$$s\bar{s}$$c\bar{c}$, or $gg$.
@@
$jj$ 4977.58 4945.68 4927.87 0.04 52.48
$ZZ\to b\bar{b}b\bar{b}$ 16.17 16.01 15.86 0.31 0.31
$ZH\to b\bar{b}b\bar{b}$ 12.14 11.88 11.70 0.31 0.31
$ZZ\to b\bar{b}jj$ 116.42 115.11 114.19 1.64 2.21
$ZH\to b\bar{b}jj$ 46.06 45.11 44.42 1.64 2.21
$ZZ\to jjjj$ 209.21 207.11 205.38 0.00 5.19
$ZH\to jjjj$ 8.60 8.42 8.29 0.00 5.19
$WW\to jjjj$ 5263.55 5202.29 5157.25 0.00 5.19
The cross section $\sigma(e^{+}e^{-}\to b\bar{b})$ is about 900 fb at around 340 GeV as shown in Fig. 4, and the $\sum_{q=u,d,s,c,b}\sigma(e^{+}e^{-}\to q\bar{q})$ is about $6000$ fb at around 340 GeV. After our event selections, around 52% of the signal events passed the event selections, while nearly ${95}\%$ of the $WW/ZZ/ZH$ background events are rejected. Particularly, more than ${99}\%$ of the four-quark backgrounds are rejected in double $b$-tagged category. Once the efficiencies of all the involved processes have been established, the $R_{b}$ values at the three energy points can be obtained together with information of integrated luminosities, the cross sections, and the estimated numbers of signal and background events.
A.4 Systematic uncertainties
The systematic uncertainties, affecting the accuracy of $R_{b}$ measurements, arise from accelerator effects (e.g., $\sqrt{s}$ energy scale and energy spread), detector efficiencies (e.g., MC statistics, $\sqrt{s^{\prime}}$ and $p_{t}$ requirement), SM parameters ($\theta_{\rm W}$, $\alpha_{s}$, etc.), theoretical considerations (e.g., higher order corrections and ISR), and others (e.g., luminosity measurements and backgrounds from two-photon and $WWZ$ processes). All the systematic uncertainties are summarized in Table 4. Taking the $R_{b}$ measurement at the second energy point as an example, systematic uncertainties are discussed below in detail.
The uncertainties related to accelerator include the $\sqrt{s}$ energy scale and energy spread. The uncertainty in the energy scale of $\sqrt{s_{2}}$ is estimated to be 2 MeV, the same as that at FCC-ee at 365 GeV [45]. The uncertainty of the energy scale contributes $14.6\times 10^{-6}$ to $\Delta R_{b}$. The energy spread is 458.4 MeV for CEPC at the second energy point. A 10% uncertainty [15, 45] in the energy spread for CEPC or FCC-ee contributes an additional $2.7\times 10^{-6}$ to $\Delta R_{b}$. The uncertainties associated with the accelerator are different at the three energy points, consistent with the $R_{b}$ distribution in Fig. 1. Notably, the first and third energy points corresponding to the minimum and maximum values of $R_{b}$ are sensitive to the energy spread, while the second energy point with the largest differential of $R_{b}$ over $m_{J_{t}}$ (very close to $\sqrt{s_{2}}$) is sensitive to the energy scale.
The uncertainty associated with the detection efficiency in $R_{b}$ is ${\varepsilon_{\text{hadron}}^{udsc}}/{\varepsilon_{\text{tag}}^{b}}$, including MC statistics and the energy-momentum requirement, and is consistent with that in the $R_{b}$ measurement at the Z pole. In addition, the sizes of simulated events can be increased to achieve a statistical error less than $10^{-4}$, thus contributing $15.4\times 10^{-6}$ to $\Delta R_{b}$. We require $p_{t}>15~{}$ GeV and $\sqrt{s^{\prime}}>\sqrt{s}-10~{}$ GeV to select signal candidates, where the resolutions of $p_{t}$ and $\sqrt{s^{\prime}}$ also affect the detection efficiency. The jet energy and angular scale of the CEPC detector can be controlled within 0.5% and 0.01% [127]. Thus, the precisions of $p_{t}$ and $\sqrt{s^{\prime}}$ can be controlled to within 0.10 GeV and 0.34 GeV, respectively, resulting in $\Delta R_{b}$ values of $9.1\times 10^{-6}$ and $6.7\times 10^{-6}$.
The dominant backgrounds originate from four-fermion final states, such as $WW$, $ZZ$, and $ZH$ related processes. These processes are well understood theoretically and experimentally, and about 95% of such events are rejected by selection criteria. After collecting large data at the $Z/WW/ZH$ peaks at CEPC and FCC-ee, and taking into account the NNLO calculation, the relative uncertainty in these backgrounds will be less than $(\alpha_{s}/\pi)^{2}\sim 0.001$, thereby contributing $13.9\times 10^{-6}$ to $\Delta R_{b}$.
The uncertainty in $\sin^{2}\theta_{\rm W}$, which can be achieved at $1\times 10^{-5}$ at CEPC [12, 45], contributes $6.0\times 10^{-6}$ to ${\Delta R_{b}}$. The uncertainty of $m_{b}$ in Eq.(15) contributes $0.2\times 10^{-6}$ to ${\Delta R_{b}}$. The uncertainty from $\alpha_{s}$ can be included in the higher order corrections. As shown in Eq.(A.2.2), the NNLO corrections will increase $R_{b}$ by $1.2\times 10^{-6}$. The uncertainty associated with ISR is canceled in $R_{b}$ measurement [43].
The other sources include luminosity measurements and backgrounds from two-photon and $WWZ$ processes. Their contributions to ${\Delta R_{b}}$ are small, with a combined estimation of $<3.0\times 10^{-6}$.
Assuming all sources of systematic uncertainties in $R_{b}$ are independent, their contributions are combined in quadrature to result in a total systematic uncertainty of $R_{b}$ at the second energy point:
$({\Delta R_{b}})_{\text{syst.}}=28.6\times 10^{-6}.$ (26)
As shown in Fig. 1, the difference in $R_{b}$ between the first and second points is $(5567.6\pm 1002.1_{\rm stat.})\times 10^{-6}$, compared to SM theoretical prediction of $-41.9\times 10^{-6}$. The total systematic uncertainty in the energy scale and energy spread is ±$\pm 39.0\times 10^{-6}$, which means that the difference in $R_{b}$ is $5.6\sigma$. Other systematic uncertainties in the detection efficiency, backgrounds, etc. do not change this difference.
@@ A.4 Systematic uncertainties
Table 4: Uncertainties of $R_{b}^{\rm PEX}$ are given in a unit of $10^{-6}$. The “Other” category includes luminosity measurements and backgrounds from two-photon and $WWZ$ processes. As a comparison, the pseudo-experiment $R_{b}^{\rm PEX}$ and SM prediction $R_{b}^{\rm SM}$ (in a unit of $10^{-6}$) are also provided here.
[…]
@@
${\rm Stat.}$ 884.4 471.2 1859.3
${\rm Total}$ 885.5 472.1 1860.4
$R_{b}^{{\rm PEX}}(\sqrt{s_{i}})$     146970.0     152537.6     156589.0
$R_{b}^{{\rm SM}}(\sqrt{s_{i}})$ 151539.9 151498.0 151466.1

### v2

@@ Appendix A End Matter
Experimental simulation—In this Letter, MC simulations of $e^{+}e^{-}$ collisions have been performed. The physics events are generated using $\tt Whizard$ [66], where the effects of ISR and energy spread have been taken into account. Subsequently, $\tt Pythia6$ [67] is applied for parton showering, hadronization, and the decay processes of the short-life hadrons. Finally, $\tt Delphes$ [68], configured with the CEPC detector conditions [13], is used to simulate the detector response. To measure $R_{b}$, samples with 2 fermions ($u\bar{u},d\bar{d},s\bar{s},c\bar{c}$,$b\bar{b}$) are generated. Moreover, due to limitations in detector performance, some jets with low energy or those produced at extreme angles relative to the detector’s central axis [13, 15] may go undetected. To address this, several multi-jet backgrounds from $WW$, $ZZ$, and $ZH$ decays to four-quark final states are simulated. In order to enhance the sensitivity for $J_{t}$ signal, the event selection criteria are optimized as discussed below.
@@
Jets from $q\bar{q}$ are clustered from all the well reconstructed tracks with an anti-kt algorithm [69]. To select high-quality jets, we require jet transverse momentum ($p_{t}$) greater than 15 GeV. To include the case of gluon final state radiation in $q\bar{q}$ final state, the selected events are supposed to have 2 or 3 jets. As the detector does not cover all directions, the ISR photons and jets close to the beam direction may escape detection. Thus, the effective invariant mass of the final states $\sqrt{s^{\prime}}$ needs to be considered. We require $\sqrt{s^{\prime}}>\sqrt{s}-10~{}{\rm GeV}$ here [70, 49, 54]. For the measurement of $R_{b}$, only events with at least two $b$-tagged jets [71] are categorized as $b\bar{b}$ events, while the remaining events are considered as light quark pair processes. For the selected signal events, we require the invariant mass of jet pairs $M_{bb}$ or $M_{j_{1}j_{2}}$ greater than 150 GeV to suppress the backgrounds from $W/Z$ hadronic decays, where ${j_{1}j_{2}}$ denote the two jets with the highest and the second-highest energies. Events passing all of the criteria are recognized as signal events, and the detection efficiencies of different processes in the three energy points are shown in Table 3.
Systematic uncertainties—The systematic uncertainties, affecting the accuracy of $R_{b}$ measurements, arise from accelerator effects (e.g., $\sqrt{s}$ energy scale and energy spread), detector efficiencies (e.g., MC statistics, $\sqrt{s^{\prime}}$ and $p_{t}$ requirements), SM parameters ($\theta_{\rm W}$, $\alpha_{s}$, etc.), theoretical considerations (e.g., higher order corrections and ISR), and others (e.g., luminosity measurements and backgrounds from two-photon and $WWZ$ processes). Taking the $R_{b}$ measurement at the second energy point as an example, systematic uncertainties are discussed below. The uncertainty in the energy scale of $\sqrt{s_{2}}$ is estimated to be 2 MeV [50], contributing $14.6\times 10^{-6}$ to $\Delta R_{b}$. The energy spread at the second energy point is 458.4 MeV for CEPC, with a 10% uncertainty [16, 50] contributing an additional $2.7\times 10^{-6}$ to $\Delta R_{b}$. By increasing the size of simulated events, the statistical error can be reduced to less than $10^{-4}$, resulting in a contribution of $15.4\times 10^{-6}$ to $\Delta R_{b}$. The precisions of $p_{t}$ and $\sqrt{s^{\prime}}$ are controlled to within 0.10 GeV and 0.34 GeV [49, 72], respectively, leading to $\Delta R_{b}$ values of $9.1\times 10^{-6}$ and $6.7\times 10^{-6}$. The $WW$, $ZZ$, and $ZH$ related processes are well understood both theoretically and experimentally, with approximately 95% of such events rejected by selection criteria. The relative uncertainty in these backgrounds is less than $(\alpha_{s}/\pi)^{2}\sim 0.001$, contributing $13.9\times 10^{-6}$ to $\Delta R_{b}$. The uncertainty in $\sin^{2}\theta_{\rm W}$, achievable at $1\times 10^{-5}$ at CEPC [13, 50], contributes $6.0\times 10^{-6}$ to $\Delta R_{b}$. The uncertainty of $m_{b}$ in Eq.(A) contributes $0.2\times 10^{-6}$ to $\Delta R_{b}$, while the uncertainty from $\alpha_{s}$ is included in higher-order corrections. As shown in Eq.(13), the NNLO corrections increase $R_{b}$ by $1.2\times 10^{-6}$. The other sources include luminosity measurements and backgrounds from two-photon and $WWZ$ processes. Their contributions to ${\Delta R_{b}}$ are small, with a combined estimation of less than $3.0\times 10^{-6}$. All the systematic uncertainties are summarized in Table 4.
@@
Table 3: The experimental cross sections $\sigma(\sqrt{s})$ (in fb), along with the detection efficiencies $\varepsilon_{\text{tag}}$ (in %, for double $b$-tagged events) and $\varepsilon_{\text{hadron}}$ (in %, for hadronic events), are provided for various processes in the $R_{b}$ measurements. Here, $jj$ denotes $u\bar{u}$, $d\bar{d}$, $s\bar{s}$, $c\bar{c}$, or $gg$.
@@
$jj$ 4,977.58 4,945.68 4,927.87 0.04 52.48
$ZZ/ZH\to b\bar{b}b\bar{b}$ $16.17/12.14$ $16.01/11.88$ $15.86/11.70$ 0.31 0.31
$ZZ/ZH\to b\bar{b}jj$ $116.42/46.06$ $115.11/45.11$ $114.19/44.42$ 1.64 2.21
$ZZ/ZH/WW\to jjjj$   $209.21/8.60/5,\!263.55$   $207.11/8.42/5,\!202.29$   $205.38/8.29/5,\!157.25$ 0.00 5.19
@@
Table 4: Uncertainties of $R_{b}^{\rm PEX}$ are given in a unit of $10^{-6}$. The “Other" category includes luminosity measurements and backgrounds from two-photon and $WWZ$ processes.
Source 340.076 GeV    341.598 GeV    342.771 GeV
@@
${\rm Stat.}$ 884.4 471.2 1,859.3
${\rm Total}$ 885.5 472.1 1,860.4